The data represent the results for a test for a certain disease. Assume one individual from the group is randomly selected. Find the probability of getting someone who tests positive​, given that he or she did not have the disease.


1
Expert's answer
2021-06-07T19:07:43-0400

Since table is not given, let us consider following example table values :

The individual actually had the disease

Yes No

Positive 121 15

Negative 39 125


"P(A/B)=\frac{P(A and B)}{P(B)}"


P(A and B)= P(test positive and do not have disease)= 15/300

P(B)= P(do not have dsease)= (15+125)/300= 140/300


Hence "P(A/B)=\frac{15}{140}=0.107"